c) Solve for internal forces just to the left of point C 5 k/ft A В C 7 3
Answer #1

I Given beam. 5K/H By equilibrium equetions { fn =0 {hyao HA=0 RA+ RB = ОВ IRA TRB (1/2x9x5) = 92.5K 18.28 3.891 EMA=0 K + MiSrn=0 5x²/18 = 4.22 → a= 3.89 ft hom D. BMX = + 4.22 (x + 1) - $x** * * BM o (n=o) = 29.54 kaft BMA (M= q) = 0 Kaft... BM e (

If you found it useful please rate the solution